# FLARE: Towards Universal Dataset Purification against Backdoor Attacks

### v1

@@ V-E Ablation Study
TABLE VI: Comparisons of FLARE (R) with SOTA backdoor-removal defenses on Tiny-ImageNet (%). The best results (highest BAs and lowest ASRs) are in bold, while BAs $<$ 50% and ASRs $>$ 20% are marked in red as fail cases.
@@
TABLE VII: Detection performance of FLARE on the CIFAR-10 dataset with different model architectures.
[…]
A2O BadNets 100.00 0.00 100.00 0.00
Blend 100.00 0.00 99.98 0.00
LC 100.00 0.00 100.00 0.00
Trojan 97.08 0.00 99.98 0.00
IAD 99.98 0.00 99.98 0.00
WaNet 94.66 0.08 95.32 0.01
ISSBA 99.98 0.00 99.90 0.00
A2A BadNets 100.00 0.00 100.00 0.00
UT BadNets 100.00 9.86 100.00 0.00
@@ V-F Resistance to Potential Adaptive Attacks
FLARE operates on the assumption that backdoored models learn distinct latent representations for poisoned and benign samples. However, studies in [13] indicate that the separations may diminish at low poisoning rates. We assess the effectiveness of FLARE against five representative attacks: BadNets (A2O), ISSBA (A2O), LC, BadNets (A2A), and BadNets (UT). These attacks are conducted on the CIFAR-10 dataset with poisoning rates ($\rho$) ranging from 0.02 to 0.1, ensuring ASRs exceed 80%. The results in Figure 8 show that FLARE maintains high effectiveness across all poisoning rates, with TPRs close to 100% and FPRs near 0%.
We further evaluate the robustness of FLARE against adaptive attacks under the worst-case scenario, where adversaries deliberately reduce the latent separation. The Ada-Patch attack [38] designs regularization samples (poisoned images with ground-truth labels) to reduce separation, posing a significant challenge for our detection. We test FLARE against Ada-Patch on the CIFAR-10 dataset with a poisoning rate of 0.1, consistent with our primary experiments. FLARE achieved a TPR of 92.10% and an FPR of 1.05%, demonstrating robustness even under this advanced adaptive attack. We argue that the effectiveness mostly stems from its ability to detect distributed anomalies across all hidden layers to accumulate subtle differences, which enhances its detection capabilities.
VI Conclusion

### v2

@@ V-E Ablation Study
TABLE VII: Comparisons of FLARE (R) with SOTA backdoor-removal defenses on Tiny-ImageNet (%). The best results (highest BAs and lowest ASRs) are in bold, while all failed cases (BA drop or ASR $>$ 10%) are marked in red.
@@
Impact of Granularity Selection in Feature Extraction. In our main experiments, FLARE extracts the minimum abnormal value at the channel level. To evaluate the impact of granularity selection on detection performance, we conduct ablation studies comparing the detection performance of layer-wise FLARE (dubbed ‘FLARE-L’) and channel-wise FLARE (dubbed ‘FLARE-C’). Specifically, FLARE-L extracts the minimum abnormal value from all channels in each layer. In contrast, FLARE-C operates at the channel level, where it extracts a minimum abnormal value from each channel separately and then aggregates these values from all channels to form the feature representation of each layer. We evaluate the performance of both FLARE-L and FLARE-C on representative backdoor attacks, including BadNets (under A2O, A2A, and UT), as well as clean-label attack (i.e., LC) and sample-specific attack (i.e., ISSBA). This experimental setup is consistent with the configurations used in other ablation studies. As shown in Table IX, FLARE-C consistently outperforms FLARE-L in terms of both TPR and FPR across all attacks. The superior performance of FLARE-C can be attributed to its ability to leverage more fine-grained feature representations, allowing it to capture more subtle differences between benign and poisoned samples, leading to better detection performance. These results verify the effectiveness of our method again.
TABLE IX: The detection performance (%) of FLARE-L and FLARE-C on the CIFAR-10 dataset.
[…]
No Poison 1.70 0.01
A2O BadNets 94.90 0.00 100.00 0.00
LC 100.00 0.00 100.00 0.00
ISSBA 88.32 0.13 99.98 0.01
A2A BadNets 89.56 0.00 100.00 0.00
UT BadNets 89.96 0.00 100.00 0.00
(a) BadNets (A2O)
(b) ISSBA (A2O)
(c) LC
(d) BadNets (A2A)
(e) BadNets (UT)
Figure 8: The impact of the poisoning rate on CIFAR-10. The consistently high TPR and low FPR across various target labels indicate stable and effective defense performance.
@@ V-F Resistance to Potential Adaptive Attacks
FLARE operates on the assumption that backdoored models learn distinct latent representations for poisoned and benign samples. However, studies in hayase2020spectre indicate that the separations may diminish at low poisoning rates. We assess the effectiveness of FLARE against five representative attacks: BadNets (A2O), ISSBA (A2O), LC, BadNets (A2A), and BadNets (UT). These attacks are conducted on the CIFAR-10 dataset with poisoning rates ($\rho$) ranging from 0.02 to 0.1, ensuring ASRs exceed 80%. The results in Figure 8 show that FLARE maintains high effectiveness across all poisoning rates, with TPRs close to 100% and FPRs near 0%.
We further evaluate the robustness of FLARE against adaptive attacks under the worst-case scenario, where adversaries deliberately reduce the latent separation. The Ada-Patch attack qi2023revisiting designs regularization samples (poisoned images with ground-truth labels) to reduce separation, posing a significant challenge for our detection. We test FLARE against Ada-Patch on the CIFAR-10 dataset with a poisoning rate of 0.1, consistent with our primary experiments. FLARE achieved a TPR of 92.10% and an FPR of 1.05%, demonstrating robustness even under this advanced adaptive attack. We argue that the effectiveness mostly stems from its ability to detect distributed anomalies across all hidden layers to accumulate subtle differences, which enhances its detection capabilities.
VI Potential Limitations and Future Work
Firstly, our method requires additional time compared to standard model training to some extent. Specifically, it involves a two-stage process: training a model on the suspicious dataset and then performing clustering on the features extracted from the model’s hidden layers to identify (potential) poisoned samples. Arguably, model training is a necessary step in all machine learning workflows, regardless of the presence of defense mechanisms. As such, the training time is not considered an additional cost. Consequently, the only additional overhead introduced by our defense is the time required for the second step. Specifically, in our experiments, training a ResNet-18 model on CIFAR-10 for 100 epochs took 1,365 seconds, whereas feature extraction and clustering took only 180 seconds (i.e., our additional overhead is 180 seconds).
Secondly, our FLARE focuses on detecting poisoned samples but does not address the reverse engineering of backdoor triggers. This task is challenging to a large extent due to the stealthiness of triggers and the lack of prior knowledge about poisoned samples. We will explore techniques to trace back and recover the triggers simultaneously in our future work.
Thirdly, our FLARE mainly focuses on image classification tasks. As multi-modal data (e.g., images, speech, and text) becomes more prevalent, we will explore extending FLARE to cross-modal scenarios to develop integrated defense strategies that combine data from multiple sources to effectively detect and prevent cross-modal backdoor attacks.
VII Conclusion